0 <t<1 3. y -5y = 10 t21 y(0)=1

Laplace transform
Answer #1

The given IvP is t² ostal 2 y-sya iyoon=1 O tol t2 ostal where f(t) = o stol - convolution theorem : :- yi-5y = f(t) 1) = {Š st žest test + - 5 25 -- viito se { **** } [) + ser]} -1) est (e-st_1) - t ist - East Bagest S tea) e 5(4-8) da s feny ese5(1-x) When t=0, S. fcase dx =0 Therefore from (t) we get Sotest when t zo y ft) = 2 + 125 est test octal > --- 125 -37 ecst-